# Memory Mosaics

### v2

@@ 4 Layered memories
In order to motivate more complex networks of memory units, let us challenge our astronomers by adding a sun and a couple planets with their respective moons. Every so often, maybe only in their imagination, our astronomers are transported to a new planetary system, observe new skies, and must sort out a new set of celestial bodies.
@@
Various hints can suggest a structure for a new planetary system: the sun has a distinctive appearance; inner planets remain close to the sun; the moons of a distant planet never stray far away from their hub, etc. Our astronomers also need geometrical knowledge to determine what to tabulate and how to combine the tabulated quantities and predict future skies. We call this knowledge persistent because it applies to all planetary systems, in contrast with contextual knowledge such as the previously observed skies for a specific planetary system.
In the three-moons architecture of Figure 3, persistent knowledge is stored in the parameters $W_{\varphi}$, $W_{\psi}$ of the feature extractors and the parameters $W_{z}$ of the mixing layer. These parameters are determined at training time through gradient back-propagation and optimization. Such simple linear functions might no longer be sufficient to encode the required persistent knowledge. Both feature extraction functions and combination functions might have to perform nonlinear computations reflecting, for instance, the projection of actual planet positions onto the celestial sphere.
Following instead the approach of Sukhbaatar et al. (2019), persistent knowledge can be stored in persistent associative memories containing a predefined number of key value pairs $(k_{i},v_{i})_{i=1\dots N_{m}}$ whose values are determined at training time through gradient back-propagation, and remain fixed at inference time. A persistent memory unit (Figure 7) is therefore very similar to a contextual memory unit (Figure 1) but relies on a persistent associative memory. Persistent memory units no longer needs an explicit value extraction function because the memory content is not updated at inference time. As pointed out in Sukhbaatar et al. (2019), they also can be viewed as fully connected neural networks with a single hidden layer that uses a soft-max non-linearity instead of a component-wise transfer function. Interleaving layers of contextual and persistent memory units can then be understood as means to increase the effective complexity of either the feature extractors or the combining layers of contextual memories (see Figure 8 for a spoiler). However we also find conceptually useful to still view the persistent memory output $y_{t}$ as the conditional expectation $\mathbb{E}(V|K)$ of an implicit value function that is not explicitly parameterized, but can be figured out after training.
@@
What kind of solutions can our astronomers design to predict the skies of a new planetary system? They may perhaps first tabulate the sun position, then the position of each planet relative to the sun, then the position of each moon relative to its planet. In order to implement such a solution, a network of memory units does not only need to compute all these relative positions, but also must route information between the memory units representing each of these tables. Furthermore, since the same network must handle sequences of observations associated with different planetary systems of different composition, these routes must change dynamically on the basis of the first observations of every new sequence of observations.
Although such dynamical routing problems seem overwhelming, we can first observe how networks of memory units can implement static routing circuits whose behavior is determined at training time. Static routing circuits are implemented using the feature extractors and the combining layer of each memory unit layer (Figure 3), both enhanced with nonlinear operations implemented by interleaved persistent memory unit layers. Just like the gating modules of a mixture of expert (Jacobs et al., 1991), such a routing circuitry can implement routes that depend on the data. However, we still call such routes static because all the possible routes are determined ahead of time during training.
In order to gain dynamic routing abilities, the training algorithm can recruit contextual memory units to replace or complement the persistent memory units involved in the routing circuitry. Because the contents of contextual memory units are updated at inference time, recruiting some of them into the routing circuitry provides the means to create new routes on the basis of the first observations of a new sequence, suggesting an alternative to capsule networks (Sabour et al., 2017).
@@
5 From planetary systems to language modeling
We have so far described Memory Mosaics as an architecture that resembles transformers in important ways and whose essential properties are comparatively transparent thanks to the predictive disentanglement principle. In order to make this vision more compelling, we provide evidence that Memory Mosaics can handle the most successful application of transformers, that is, language modeling.
Let us first argue that predictive disentanglement is a meaningful concept for language modeling. In the case of a dialogue, predicting who is going to speak next can largely be predicted independently from what is being said. The same applies to predicting whether the next piece of dialog is expressed in English or German, whether the participants are in agreement or not, or whether their grammar is formal or informal. Disentangling such predictions provide opportunities to leverage more training data. For instance, one can predict that a specific dialog participant will provide a tense answer because virtually any dialog participant placed the same a situation would provide a tense answer.
@@ 5 From planetary systems to language modeling
The TinyStories dataset of Eldan and Li (2023) provides the means to use relatively small language models to study language modeling phenomena that usually only appear in much larger models. This is achieved by constraining the text written in the simple language that a three years old child could understand and recounting stories that take place in the simple world that a three years old child could understand. Thanks to this limited scope, Eldan and Li show how a small language model (33M parameters) trained on a corpus of such tiny stories generates continuations with far better language quality and narrative consistency than those of a much larger model (1.5B parameters) trained on a generic web corpus.
Following both the lead of Eldan and Li and the admonitions of our legal department, we use the Mixtral-8x7B open language model (Jiang et al., 2024) to generate a new corpus of tiny stories dubbed BabiStories. This corpus and its generation are detailed in Appendix A.
@@
To put our experiments into context, we design a Memory Mosaic architecture that closely matches the classic GPT2-small transformer architecture (Radford et al., 2018, 2019). Both architectures, shown side-by-side in Figure 8, use the same GPT2 tokenizer, the same embedding dimension ($d=768$), and the same number of heads ($N_{h}=N_{c}=N_{p}=12$). Both architectures are trained and tested using sequences of length 512, that is, one to three stories long.
There are three major differences between these two architectures. First, the Memory Mosaic does not use positional encoding. Second, unlike the $N_{h}=12$ attention heads of each transformer block, the $N_{c}=12$ contextual memory units in each block do not distinguish keys from queries (Figure 1) but instead use the key and value extraction functions described in Equation 7. The keys are formed with a leaky average of past inputs, and the values can peek one time step ahead. Accordingly, the attention mask excludes the main diagonal to avoid breaking causality. Finally, the feed forward networks (FFNs) of the classic transformers blocks are replaced by a layer of $N_{p}=12$ persistent memory units, complete with a key extraction functions (7) and combining layer, sized to ensure that the per-block parameter count of the Memory Mosaic architecture closely matches GPT2-small. [^4] Compared with GPT2-small, we save ${768\times 512}$ position encoding weights and ${N_{b}\times 768^{2}}$ query projection weights, and add ${2\times{N_{b}}\times 768^{2}}$ weights for the persistent memory key extraction function and mixing layer. The total number of persistent memory unit slots is therefore close, but not exactly equal, to the number of FFN hidden units.
Figure 9: Training and validation loss of the transformer and Memory Mosaic architectures trained on BabiStories for different model depths. The horizontal axis represents the number of training iterations. All hyper-parameters have been tuned on the transformer architecture and transferred verbatim to the Memory Mosaic architecture. The Memory Mosaic slightly outperforms the transformer for small depth networks, but that effect disappears when the depth increases.
@@
Figure 9 shows the training and validation curves of both transformers and Memory Mosaics of different depth trained on BabiStories. The Memory Mosaic slightly outperforms the transformer for small depth networks, [^5] This is not surprising because the Memory Mosaic architecture only needs one block to correctly handle the induction head problem of Bietti et al. (2024), whereas the transformer architecture needs two blocks for the same task., but this effect disappears when the depth increases and both the training and validation losses become indistinguishable.
@@
Importantly, all hyper-parameters were tuned for the transformer architectures (Appendix B) and transferred verbatim to the Memory Mosaics. This choice might explain why the training curves track each other so well. It also leaves the Memory Mosaics at a slight disadvantage in this comparison.
@@
Table 1: Continuations generated for the 24 prompts designed by Eldan and Li (2023) to investigate the factual, reasoning, and consistency capabilities of language models trained on BabiStories. Both the transformer and the Memory Mosaic were $N_{b}=18$ blocks deep. Colors green and red respectively indicate correct and incorrect answers. The text was generated with a beam search of width 2 and manually truncated to the shortest length that answers or fails to answer the question.
@@
Qualitative evaluation
In order to compare the quality of the text generated by models trained on tiny stories, Eldan and Li designed twenty-four prompts that exercise the factual, logical, and consistency properties of the generated continuations. Table 1 compares the continuation generated on these prompts by a transformer and a Memory Mosaic, both $N_{b}=18$ blocks deep. Both models perform very similarly on this task.
Figure 11: Memory Mosaics performance on the RegBench in-context learning benchmark (Akyürek et al., 2024). Since RegBench includes a search over architectural hyper-parameters, Memory Mosaics and transformers use the same search space with the same parameter counts. Memory mosaics outperform all previously tested architectures in this benchmark.
Out-of-distribution evaluation
The Simple English Wikipedia [^6] https://simple.wikipedia.org/wiki/Simple_English_Wikipedia is a version of Wikipedia written in a language that is easier to understand. Despite the intended simplicity, the articles are substantially longer and more sophisticated than our BabiStories. Predicting Simple English Wikipedia articles using models trained on BabiStoriesis therefore a challenging out-of-distribution task.
Figure 10 shows the per-token average loss as a function of the position of the generated token in the input window. Both the transformer and the Memory Mosaic are $N_{b}=512$ blocks deep. In this experiment, the token prediction is expected to improve when the increasing context size reveals that the distribution is different. The transformer performance plateaus after 100 to 150 tokens, which is a bit shorter than a typical tiny story. Memory Mosaics substantially outperform transformers after about 50 tokens, suggesting superior in-context learning abilities.
In-context learning evaluation
In order to rigorously compare the in-context learning abilities of various architectures, the RegBench benchmark (Akyürek et al., 2024) constructs a large number of artificial languages defined by probabilistic finite automatons (PFA) constructed at random. Each input sequence is composed of 10 to 20 strings drawn from a same PFA and delimited separator tokens. The competing architectures are trained on a variable number of input sequences, then evaluated on their ability to predict the last token of testing sequences generated using held out PFAs.
Since RegBench performs a hyper-parameter search for each data point, we use the Memory Mosaic architecture of Figure 8 with the same search space as transformers, ensuring that both transformers and Memory Mosaics have the same parameter count for the same architectural hyper-parameters. We sweep over depth $N_{b}\in\{2,4,8\}$, number of heads $N_{h}{=}N_{c}{=}N_{p}\in\{2,4,8\}$, embedding dimension in $d\in\{64,128,256\}$, weight decay in $\{10^{-2},10^{-1}\}$, and training epochs in $\{1,2,\dots 200\}$.
Figure 11 compares Memory Mosaic on RegBench with the results previously reported by Akyürek et al.. The left plot shows the prediction accuracy for the test string last token. The right plot compares the predicted last token distribution with the exact distribution implied by PFA. Memory Mosaics dominate this benchmark, substantially outperforming transformers, recurrent neural networks, and state-space models for training set sizes covering three orders of magnitude. [^7] In the low train environments case (e.g. 100), all baseline methods which perform poorly in the RegBench OOD test set can actually perform very well in a IID test set. Thus baseline methods learned the training environments (good IID) but not the meta-learning ability (poor OOD). Please check appendix table 4 for details.
Attention differences
Because Memory Mosaics lack position encoding and do not distinguish keys and queries, we investigate how their attention patterns differ from those of transformers. Figure 12 shows attention scores for each head of either a one-block deep transformer using absolute position encoding (left plot) or a one-block deep Memory Mosaic (right plot). The scores are averaged on 5000 BabiStories sequences and show how the last position attends to earlier positions in the 512 token long context window. The transformer attention patterns are noisy, with a strong “attention sink” at position 0 (Xiao et al., 2023). In contrast, the Memory Mosaic attention pattern is mostly flat, save for higher scores for the most recent tokens. [^8] This effect is connected to the leaky average coefficient $\lambda_{\varphi}$, as shown in Appendix (Figure 15).
6 Discussion
The starting point of this work is made of two very old ideas. The first one is augment a deep network with explicit memories. The second one is to let the learning process decide what gets memorized and how it gets retrieved. Although such ideas have been explored in memory networks (Weston et al., 2014; Joulin and Mikolov, 2015; Sukhbaatar et al., 2015), the importance of having lots of independent memories had not been fully appreciated.
This contribution focuses on networks of associative memories implemented with kernel smoothing, therefore amenable to gradient-based learning algorithms. Such learning machines not only resemble decoding transformers (Section 2) but also perform very much like decoding transformers on the sort of language modeling task that made them famous (Section 5). Although much work is needed to replicate our observations at far greater scale, Memory Mosaics satisfy narrative constraints as well as transformers (Table 1), and generally behave in very encouraging ways (Figures 10 to 13).
Memory Mosaics also offer an array of engineering opportunities. Limited storage contextual memories could be implemented by evicting the least-recently used entry rather than by bounding the context size (see also (Xiao et al., 2023)). Associative memories can be implemented using a wide spectrum of techniques, ranging from fast transforms (Greengard and Strain, 1991; Yang et al., 2003) to locality sensitive hashing mechanisms (Spring and Shrivastava, 2017; Chen et al., 2020) that could change the computing requirements of contemporary artificial intelligence systems.
Acknowledgements
JZ and LB received partial funding from the CIFAR Learning in Machines & Brains program.
Appendix A BabiStories
The TinyStories dataset [Eldan and Li, 2023] is composed of stories written in a simple language and taking place a narrow world. Such stories can be used to train relatively small language models that still must address some of the broader language modeling challenges such as obeying narrative necessity and maintaining logical consistency. This dataset is a wonderful way to study big problems with acceptable computation and quick turn around.
The experiments of Section 5 were carried out using a dataset generated using a similar methodology but using the Mixtral-8x7B open language model in order to generate unencumbered data. We call this dataset BabiStories. All the scientific credit is still due to the remarkable work of Eldan and Li. Table 2 provides basic statistics for this newly generated BabiStoriesdataset, essentially matching those of the original TinyStories dataset of Eldan and Li [2023]. We had to increase the diversity of the generated stories by expanding the prompt to specify first names and by providing opening words for the story, in addition to required words and story features used by Eldan and Li (Figure 14). We also removed the few generated stories containing URLs.
Figure 14: Generation of BabiStories. In order to improve the diversity of the generations, each story is generated by a prompt that provides a list of required words and story features (as in Eldan and Li, 2023) and additionally provides first names and opening words.
Table 2: BabiStories statistics.
train 2.2M 474,704,907 888
valid 2.2k 4,749,107 889
Appendix B GPT2 baseline and hyperparameters
Table 3 showcases the hyper-parameters searching process of GPT2 transformer baseline on the BabiStories dataset, where we use AdamW optimizer Loshchilov and Hutter [2017], batch-size 512, context-size 512, and a cosine learning rate scheduler with minimum learning rate $41e-$ 4 for all training.
5e-3 0.05 0.1 2000 80000 1.336 1.494
1e-3 0.05 0.1 2000 80000 1.350 1.524
5e-3 0 0.1 2000 80000 1.281 1.556
5e-3 0.05 0.01 2000 80000 1.322 1.516
5e-3 0.05 0.1 200 80000 fail fail
5e-3 0.05 0.1 2000 40000 1.325 1.532
5e-3 0.05 0.1 2000 160000 1.314 1.497
Appendix C Memory Mosaics for language modeling
C.1 Additional qualitative evaluation
Table 5 provides a similar continuation generation comparison as Table 1 in Section 5, but setting $N_{b}=1$.
C.2 Attention map and leaky average coefficient $\lambda_{\varphi}$
C.3 In-context language learning
Table 4 provides the IID test performance of various architectures trained on RegBench [Akyürek et al., 2024] with 100 training environments. We keep the training process, including hyperparameter searching space, to be the same as the one in Figure 11. But sample validation and test sets from the same 100 probabilistic finite automatons (training environments) as the training set. This table, together with Figure 11, show that baseline methods learned the training environments (good IID) but not the meta-learning ability (poor OOD).
Memory Mosaic tf Mamba S4 RWKV linear tf H3 GLA Hyena LSTM RetNet
Table 4: In-distribution (IID) performance of various architectures trained on RegBench [Akyürek et al., 2024] with only 100 training environments. Both training, validation, and test set (100 samples) are sampled from the same 100 random probabilistic finite automatons (PFA). Compared with the poor OOD accuracy ($\sim$0.45) / TVD ($\sim$0.75) of baseline methods in Figure 11, All baseline methods perform well in the IID test set (even with only 100 training environments).
@@ C.3 In-context language learning
Table 5: Continuations generated for the 24 prompts designed by Eldan and Li [2023] to investigate the factual, reasoning, and consistency capabilities of language models trained on BabiStories. Both the transformer and the Memory Mosaic were $N_{b}=1$ blocks deep. Colors green and red respectively indicate correct and incorrect answers. The text was generated with a beam search of width 2 and manually truncated to the shortest length that answers or fails to answer the question.

### v3

@@ 6 Layered memories
We of course envision deeper networks of memory units. In order to make meaningful comparisons, we also would like to remain as close as possible to the classic transformer architecture which alternates self-attention layers with fully connected feed-forward networks (FFNs).
@@
Sukhbaatar et al. (2019) shows that FFNs in a transformer can be interpreted as persistent memories that augment the self-attention layers and provide means to represent information that persists across input sequences. Besides the contextual memory units (Figure 1), we therefore introduce persistent memory units (Figure 12 in the Appendix) that contain a predefined number of key value pairs $(k_{i},v_{i})_{i=1\dots N_{m}}$ determined at training time through gradient back-propagation. Persistent memory units no longer need an explicit value extraction function because the memory content is not updated at inference time. As pointed out by Sukhbaatar et al., they also can be viewed as fully connected neural networks with a single hidden layer that uses a soft-max non-linearity instead of a component-wise transfer function. Yet, we find conceptually useful to still view the persistent memory output $y_{t}$ as the conditional expectation $\mathbb{E}(V|K)$ of an implicit value function that is not explicitly parameterized, but can be figured out after training.
@@
Interleaving layers of contextual and persistent memory units can then be understood as means to increase the effective complexity of either the feature extractors or the combining layers of contextual memories (see Figure 6 for a spoiler). Therefore persistent memory units can also be seen as tool for routing information between successive layers of contextual memory units. Such a circuitry can implement routes that depend on the data, just like the gating modules of a mixture of expert (Jacobs et al., 1991). Since all the parameters of such a circuitry are determined at training time, all the possible routes would have to be determined at training time. However the learning algorithm can overcome this limitation by also recruiting contextual memory units from adjacent layersy. Because the contents of contextual memory units are updated at inference time, recruiting some of them into the routing circuitry provides the means to create new routes on the basis of the first observations of a new sequence, suggesting an efficient alternative to capsule networks (Sabour et al., 2017).
@@
7 Modeling language with memories
We have so far described Memory Mosaics as an architecture that resembles transformers in important way but offers additional insights such as predictive disentanglement. We now provide evidence that Memory Mosaics can handle the most successful application of decoding transformers, that is, language modeling.
@@ 7 Modeling language with memories
The TinyStories work of Eldan & Li (2023) shows how to study large language modeling questions using small language models. This is achieved by limiting the scope to tiny stories written in simple english and taking place in the simple world that a three years old child could understand. A small language model trained on such data generates continuations with far better language quality and narrative consistency than those a much larger model (1.5B parameters) trained on a generic text.
Following both the lead of Eldan & Li and the advice of our legal department, we leverage the Mixtral-8x7B open language model (Jiang et al., 2024) to generate a new corpus of tiny stories dubbed BabiStories. This corpus and its generation are detailed in Appendix B. [^5] We share the BabiStories dataset and Memory Mosaics source code at https://github.com/facebookresearch/MemoryMosaics.
@@
To put our experiments into context, we design a Memory Mosaic architecture that closely matches the classic GPT2-small transformer architecture (Radford et al., 2018; 2019). Both architectures, shown side-by-side in Figure 6, use the same GPT2 tokenizer, the same embedding dimension ($d=768$), and the same number of heads ($N_{h}=N_{c}=N_{p}=12$}). Both architectures are trained and tested using sequences of length 512, that is, one to three stories long.
There are three major differences between these two architectures. First, the Memory Mosaic does not use positional encoding. Second, unlike the $N_{h}=12$ attention heads of each transformer block, the $N_{c}=12$ contextual memory units in each block do not distinguish keys from queries (Figure 1) but instead use the key and value extraction functions described in Equation 6. The keys are formed with a leaky average of past inputs, and the values can peek one time step ahead. [^6] The key idea here is to define key and value extraction functions that combine a couple successive inputs $x_{t}$ instead of just one as in the three moons example. Many variations perform more or less equivalently. Accordingly, the attention mask excludes the main diagonal to avoid breaking causality. Finally, the feed forward networks (FFNs) of the classic transformers blocks are replaced by a layer of $N_{p}=12$ persistent memory units, complete with a key extraction functions (6) and combining layer. These persistent memory units are sized to ensure that the per-block parameter count of the Memory Mosaic architecture closely matches GPT2-small. [^7] Compared with GPT2-small, we save $768\times 512$} position encoding weights and $N_{b}\times 768^{2}$} query projection weights, but add $2\times{N_{b}}\times 768^{2}$} weights for the persistent memory key extraction and mixing layer. The total number of persistent memory unit slots is therefore close to the total number of FFN hidden units.
@@
Figure 7: Training and validation loss of the transformer and Memory Mosaic architectures trained on BabiStories for different model depths. The horizontal axis represents the number of training iterations. All hyper-parameters have been tuned on the transformer architecture and transferred verbatim to the Memory Mosaic architecture. The Memory Mosaic slightly outperforms the transformer for small depth networks, but that effect disappears when the depth increases. Additional results are presented in Appendix D.2.
@@
Figure 7 shows the training and validation curves of both transformers and Memory Mosaics of different depth trained on BabiStories. The Memory Mosaic slightly outperforms the transformer for small depth networks, [^8] This is not surprising because Memory Mosaics only need a single block to implement induction heads, whereas transformers need at least two for the same task. but this effect disappears when the depth increases and both the training and validation losses become indistinguishable. Additional results are presented in Appendix D.2.
@@
Importantly, all hyper-parameters were tuned for the transformer architectures (Appendix C) and transferred verbatim to the Memory Mosaics. This choice might explain why the training curves track each other so well. It also leaves the Memory Mosaics at a slight disadvantage.
Qualitative evaluation
In order to compare the quality of the text generated by models trained on tiny stories, Eldan & Li designed twenty-four prompts that exercise the factual, logical, and consistency properties of the generated continuations. Table 4 in the Appendix compares the continuation generated on these prompts by a transformer and a Memory Mosaic, both $N_{b}=18$ blocks deep. Both models perform very similarly on this task.
[Uncaptioned image] [Uncaptioned image]
Figure 9: Memory Mosaics performance on the RegBench in-context learning benchmark (Akyürek et al., 2024). Since RegBench includes an hyper-parameter search, Memory Mosaics and transformers use the same search space with the same parameter counts. Memory mosaics outperform all previously tested architectures in this benchmark.
Out-of-distribution evaluation
The Simple English Wikipedia [^9] Described in https://simple.wikipedia.org/wiki/Simple_English_Wikipedia with downloads in https://huggingface.co/datasets/wikipedia#20220301simple. is a version of Wikipedia written in a language that is easier to understand. Despite the intended simplicity, the articles are substantially longer and more sophisticated than our BabiStories. Predicting Simple English Wikipedia articles using models trained on BabiStoriesis therefore a challenging out-of-distribution task.
Figure 9 shows the per-token average loss as a function of the position of the generated token in the input window. Both the transformer and the Memory Mosaic are $N_{b}=12$ blocks deep. In this experiment, the token prediction is expected to improve when the increasing context size reveals that the distribution is different. The transformer performance plateaus after 100 to 150 tokens, which is a bit shorter than a typical tiny story. Memory Mosaics substantially outperform transformers after about 50 tokens, suggesting superior in-context learning abilities.
In-context learning evaluation
In order to rigorously compare the in-context learning abilities of various architectures, the RegBench benchmark (Akyürek et al., 2024) constructs random artificial languages defined by probabilistic finite automata (PFA). Each input sequence is composed of 10 to 20 strings drawn from a same PFA and delimited separator tokens. The competing architectures are trained on a variable number of input sequences, then evaluated on their ability to predict the last token of testing sequences generated using held out PFAs.
Since RegBench performs a hyper-parameter searches, we use the Memory Mosaic architecture of Figure 6 with the same search space as transformers, ensuring that both transformers and Memory Mosaics have the same parameter count for the same architectural hyper-parameters. We sweep over depth $N_{b}\in\{2,4,8\}$, number of heads $N_{h}{=}N_{c}{=}N_{p}\in\{2,4,8\}$, embedding dimension in $d\in\{64,128,256\}$, weight decay in $\{10^{-2},10^{-1}\}$, and training epochs in $\{1,2,\dots 200\}$.
Figure 9 compares Memory Mosaic on RegBench with the results previously reported by Akyürek et al.. The left plot shows the prediction accuracy for the test string last token. The right plot compares the predicted last token distribution with the exact distribution implied by PFA. Memory Mosaics dominate this benchmark, substantially outperforming transformers, recurrent neural networks, and state-space models for training set sizes covering three orders of magnitude. [^10] Although the baseline methods trained with small training sets (e.g. 100) perform poorly on the RegBench task, they perform very well when tested in-distribition (see Table 3 in the Appendix). Therefore they learned to model the training languages but did not acquire the ability to learn new languages in context.
8 Discussion
The starting point of this work is made of two very old ideas. The first one is augment a deep network with explicit memories. The second one is to let the learning process decide what gets memorized and how it gets retrieved. Although such ideas have been explored in memory networks (Weston et al., 2014; Joulin & Mikolov, 2015; Sukhbaatar et al., 2015), the importance of having lots of independent memories had not been fully appreciated.
This contribution focuses on networks of associative memories implemented with kernel smoothing, therefore amenable to gradient-based learning algorithms. Such learning machines not only resemble decoding transformers (Section 3) but also perform very much like decoding transformers on the sort of language modeling task that made them famous (Section 7). Although much work is needed to replicate our observations at far greater scale, Memory Mosaics satisfy narrative constraints as well as transformers (Table 4), and generally behave in very encouraging ways (Figures 9 to 15).
Memory Mosaics also offer an array of engineering opportunities. Limited storage contextual memories could leverage least-recently used eviction schemes (e.g., Xiao et al., 2023), and associative memories could be implemented using a wide spectrum of techniques, either classical (e.g., Greengard & Strain, 1991; Spring & Shrivastava, 2017), or neural (e.g., Krotov, 2023), which could redefine the computing requirements of contemporary artificial intelligence systems.
Appendix A Tracking three moons
Figure 10 shows how the training process yields parameter matrices $W_{\varphi}$, $W_{\psi}$, and $W_{z}$, that dedicate one memory unit to each moon.
Appendix B BabiStories
The TinyStories dataset (Eldan & Li, 2023) is composed of stories written in a simple language and taking place a narrow world. Such stories can be used to train relatively small language models that still must address some of the broader language modeling challenges such as obeying narrative necessity and maintaining logical consistency. This dataset is a wonderful way to study big problems with acceptable computation and quick turn around.
The experiments of Section 7 were carried out using a dataset generated using a similar methodology but using the Mixtral-8x7B open language model in order to generate unencumbered data. We call this dataset BabiStories. All the scientific credit is still due to the remarkable work of Eldan & Li. Table 1 provides basic statistics for this newly generated BabiStoriesdataset, essentially matching those of the original TinyStories dataset of Eldan & Li (2023). We had to increase the diversity of the generated stories by expanding the prompt to specify first names and by providing opening words for the story, in addition to required words and story features used by Eldan & Li (Figure 11). We also removed the few generated stories containing URLs.
Figure 11: Generation of BabiStories. In order to improve the diversity of the generations, each story is generated by a prompt that provides a list of required words and story features (as in Eldan & Li, 2023) and additionally provides first names and opening words.
Table 1: BabiStories statistics.
train 2.2M 474,704,907 888
valid 2.2k 4,749,107 889
Appendix C GPT2 baseline and hyperparameters
Table 2 showcases the hyper-parameters searching process of GPT2 transformer baseline on the BabiStories dataset, where we use AdamW optimizer Loshchilov & Hutter (2017), batch-size 512, context-size 512, and a cosine learning rate scheduler with minimum learning rate 1e41e-4 for all training.
5e-3 0.05 0.1 2000 80000 1.336 1.494
1e-3 0.05 0.1 2000 80000 1.350 1.524
5e-3 0 0.1 2000 80000 1.281 1.556
5e-3 0.05 0.01 2000 80000 1.322 1.516
5e-3 0.05 0.1 200 80000 fail fail
5e-3 0.05 0.1 2000 40000 1.325 1.532
5e-3 0.05 0.1 2000 160000 1.314 1.497
Appendix D Memory Mosaics for language modeling
D.1 Persistent memory units
Persistent memory units produce their outputs using the same key extraction function $\varphi(x_{T},x_{T-1},\dots)$ and the same retrieval function (3) as contextual memory units. They differ because, following Sukhbaatar et al. (2019), they use a fixed array of key/values pairs that are treated as parameters and are determined at training time by gradient descent. Since these stored key/value pairs do not change at inference time, there is no need for a value extraction function $\psi(x_{T+1},x_{T},\dots)$
D.2 Training and validation
Figure 13 plots the training and validation curves for both Transformer and Memory Mosaic in a manner similar to Figure 7 but showing additional block depths.
Figure 13: Additional training and validation curves for the transformer and Memory Mosaic architectures trained on BabiStories for more model depths than Figure 7.
Several comments can be made:
  • The Memory Mosaic has a small advantage for very small depths ($N_{b}=1$ and $N_{b}=4$) but this advantage does not persist when the number of blocks increases. We believe this is due to the fact that a single layer Memory Mosaic can implement an induction head whereas a Transformer needs two layers. This amounts to saying that a $n$ block deep Mosaic has the same number of parameters than a $n$ block deep Transformer, its performance is closer to that of a $n+1$ block Transformer. This is not much of an advantage when $n$ gets large.
  • The Memory Mosaic training uses the hyper-parameters that worked best for the Transformer and operates on the same mini-batches of examples in the same order. However, for small block depths, the Memory Mosaic training curve shows initial instability, suggesting that it might benefit from a smaller stepsize.
  • The similarity of the Transformer and Memory Mosaic curves is especially striking when one recalls that the Memory Mosaic does not use position encoding. In fact Memory Mosaic have two mechanisms for dealing with positions. The first one is the fact that the values $v_{T}$ peek one time position ahead. The second one is the leaky integration in (6). These two mechanisms are useful to implement bigram or n-gram induction heads in a single layer, but they do not allow a head to selectively address a token by position (we use a single scalar leaky average coefficient per head). This suggests that position encoding in Transformers is mostly useful to implement an initial induction head in the first two blocks.
D.3 Qualitative evaluation
Table 5 provides a variant of Table 4 in Section 7, with $N_{b}=1$.
D.4 Differences in Attention and the leaky average coefficient $\lambda_{\varphi}$
Because Memory Mosaics lack position encoding and do not distinguish keys and queries, we investigate how their attention patterns differ from those of transformers. Figure 14 shows attention scores for each head of either a one-block deep transformer using absolute position encoding (left plot) or a one-block deep Memory Mosaic (right plot). The scores are averaged on 5000 BabiStories sequences and show how the last position attends to earlier positions in the 512 token long context window. The transformer attention patterns are noisy, with a strong “attention sink” at position 0 (Xiao et al., 2023). In contrast, the Memory Mosaic attention pattern is mostly flat, save for higher scores for the most recent tokens. [^13] This effect is connected to the leaky average coefficient $\lambda_{\varphi}$, as shown in Figure 16.
D.5 In-context language learning evaluation
Table 3 provides the IID test performance of various architectures trained on RegBench (Akyürek et al., 2024) with 100 training environments. We keep the training process, including hyperparameter searching space, to be the same as the one in Figure 9. But sample validation and test sets from the same 100 probabilistic finite automatons (training environments) as the training set. This table, together with Figure 9, show that baseline methods learned the training environments (good IID) but not the meta-learning ability (poor OOD).
Table 3: In-distribution (IID) performance of various architectures trained on RegBench (Akyürek et al., 2024) with only 100 training environments. Both training, validation, and test set (100 samples) are sampled from the same 100 random probabilistic finite automatons (PFA). Compared with the poor OOD accuracy ($\sim$0.45) / TVD ($\sim$0.75) of baseline methods in Figure 9, All baseline methods perform well in the IID test set (even with only 100 training environments).
Memory Mosaic tf Mamba S4 RWKV linear tf H3 GLA Hyena LSTM RetNet
Appendix E Computing Resources
Models were trained on 64 NVidia V100 GPUs over 80k epochs. From conception to finalization of this paper we trained about 200 models. To create the BabiStories dataset via Mistral, we ran with 128 NVidia V100 GPUs for 3 days. The supporting machines contain Intel(R) Xeon(R) Gold 6230 CPUs. The 3 moons result took negligible resources and were trained on Apple M1 laptops.
@@ Appendix E Computing Resources
Table 4: Continuations generated for the 24 prompts designed by Eldan & Li (2023) to investigate the factual, reasoning, and consistency capabilities of language models trained on BabiStories. Both the transformer and the Memory Mosaic were $N_{b}=18$ blocks deep. Colors green and red respectively indicate correct and incorrect answers. The text was generated with a beam search of width 2 and manually truncated to the shortest length that answers or fails to answer the question.
@@
Table 5: Continuations generated for the 24 prompts designed by Eldan & Li (2023) to investigate the factual, reasoning, and consistency capabilities of language models trained on BabiStories. Both the transformer and the Memory Mosaic were $N_{b}=1$ blocks deep.